what is 41 3/4 percent as a fraction and also what is 61 1/4 percent as a fraction
Answer by tommyt3rd(5050)

41 3/4 percent means
41.75/100 =
4175/1000=
(167*25)/(400*25)=
167/400

61 1/4 percent means
6125/1000=
(49*125)/(80*125)=
49/80